OpenStudy (anonymous):

1/b divided by 1/a-1/b anyone wanna try take a whack at that?

OpenStudy (anonymous):

First need to combine the 1/a - 1/b. Multiply the first by b/b and the second by a/a to make them have a common denominator. b/ab - a/ab = (b-a)/ab Now we can use the reciprocal of that to do the division. 1/b * [(ab)/(b-a)] You can multiply and simplify from there.

OpenStudy (anonymous):

little confused

OpenStudy (anonymous):

Okay. If I have 1/4 - 1/5 I need to make a common denominator in order to subtract them. In this case I would multiply the first by 5/5 and the second by 4/4 to make them have the same denominator of 20. The same applies when you have 1/a - 1/b only this time with variables. I multiply the first by b/b and the second by a/a so they will both have a denominator of ab.

OpenStudy (anonymous):

(1/a) * (b/b) = b/ab (1/b) * (a/a) = a/ab b/ab - a/ab = (b-a)/ab

OpenStudy (anonymous):

Then I need to take 1/b divided by (b-a)/ab So I keep the first fraction, change the operation to multiplication, and flip the second fraction (this is how you multiply fractions). (1/b) * ab/(b-a) Then you can just multiply across the numerator and denominator like you multiply any two fractions. The numerator is ab. The denominator is b(b-a)